# Zero-Order Sharpness-Aware Minimization

### v1

@@ A.4.1 SAM Mechanism
A.4.2 ZOSA Approximation
ZOSA operates in a zero-order setting, meaning it does not have access to exact gradients. Instead, it estimates gradients using function evaluations along random directions. The step-by-step derivation of how ZOSA simulaties SAM is shown below.
@@ A.4.2 ZOSA Approximation
    $\epsilon_{\text{sam}}\approx\rho\frac{\nabla L(\theta_{t})}{\|\nabla L(\theta_{t})\|}.$ (66)
@@ A.4.3 Conclusion for Equivalence
The proof builds on standard non-convex descent lemmas, ZO estimation bias/variance bounds, SAM perturbation approximations, and Property 3.2 ($\sigma_{t}\approx\epsilon\|\nabla\|$).
Assume $L$ is $\mathcal{L}$-smooth and non-convex with lower bound $L^{*}$, perturbations satisfy $\mathbb{E}[(l_{i}-L(\theta))^{2}]\leq\sigma^{2}$, $\|\nabla L\|\leq G$, $\eta\leq\frac{m}{16d\mathcal{L}}$, $\epsilon\leq\frac{1}{\sqrt{d\mathcal{L}}}$, $\rho\leq\frac{1}{4\mathcal{L}}$. By $\mathcal{L}$-smoothness:
@@
where $\Delta\theta=-\frac{\eta}{\sigma_{t,\text{pert}}}g_{\text{pert}}$. Substitute:
$L(\theta_{t+1})\leq L(\theta_{t})-\frac{\eta}{\sigma_{t,\text{pert}}}\langle\nabla L(\theta_{t}),g_{\text{pert}}\rangle+\frac{\mathcal{L}}{2}\left(\frac{\eta}{\sigma_{t,\text{pert}}}\right)^{2}\|g_{\text{pert}}\|^{2}.$ (73)
@@
$\mathbb{E}[L(\theta_{t+1})]\leq L(\theta_{t})-\frac{\eta}{\sigma_{t,\text{pert}}}\langle\nabla L(\theta_{t}),\mathbb{E}[g_{\text{pert}}]\rangle+\frac{\mathcal{L}}{2}\left(\frac{\eta}{\sigma_{t,\text{pert}}}\right)^{2}\mathbb{E}[\|g_{\text{pert}}\|^{2}].$ (74)
From ZO properties at perturbed point $\theta_{\text{pert}}=\theta_{t}+\rho\frac{g_{t}}{\sigma_{t}}$:
$\mathbb{E}[g_{\text{pert}}]=\nabla L(\theta_{\text{pert}})+O(\epsilon^{2}d\mathcal{L}),$ (75)
$\mathbb{E}[\|g_{\text{pert}}-\mathbb{E}[g_{\text{pert}}]\|^{2}]\leq\frac{d(\sigma^{2}+G^{2}\epsilon^{2})}{\epsilon^{2}m},$ (76)
$\mathbb{E}[\|g_{\text{pert}}\|^{2}]=\|\nabla L(\theta_{\text{pert}})\|^{2}+O\left(\frac{d(\sigma^{2}+G^{2}\epsilon^{2})}{\epsilon^{2}m}\right).$ (77)
By smoothness, $\|\nabla L(\theta_{\text{pert}})-\nabla L(\theta_{t})\|\leq\mathcal{L}\rho$, since $\left\|\frac{g_{t}}{\sigma_{t}}\right\|\approx 1$ (from Property 3.2, $\sigma_{t}\approx\epsilon\|\nabla L(\theta_{t})\|$, and $g_{t}\approx\nabla L(\theta_{t})$ up to scaling adjustment for one-sided Rademacher). Thus:
$\langle\nabla L(\theta_{t}),\nabla L(\theta_{\text{pert}})\rangle$ $=\langle\nabla L(\theta_{t}),\nabla L(\theta_{t})+(\nabla L(\theta_{\text{pert}})-\nabla L(\theta_{t}))\rangle$ (78)
$\geq\|\nabla L(\theta_{t})\|^{2}-\mathcal{L}\rho\|\nabla L(\theta_{t})\|,$ (79)
$\langle\nabla L(\theta_{t}),\mathbb{E}[g_{\text{pert}}]\rangle$ $\geq\|\nabla L(\theta_{t})\|^{2}-\mathcal{L}\rho\|\nabla L(\theta_{t})\|$ (80)
$+O(\epsilon^{2}d\mathcal{L}\|\nabla L(\theta_{t})\|).$ (81)
For bounding, assume $\eta/\sigma_{t,\text{pert}}\leq\eta/(\epsilon\sqrt{\delta})$ for small $\delta>0$ (handling near-zero division), and $\|\nabla L(\theta_{\text{pert}})\|^{2}\leq 2\|\nabla L(\theta_{t})\|^{2}+2(\mathcal{L}\rho)^{2}$:
$-\frac{\eta}{\sigma_{t,\text{pert}}}\langle\nabla L(\theta_{t}),\mathbb{E}[g_{\text{pert}}]\rangle$ $\leq-\frac{\eta}{\sigma_{t,\text{pert}}}\left(\frac{1}{2}\|\nabla L(\theta_{t})\|^{2}-\rho\mathcal{L}\|\nabla L(\theta_{t})\|^{2}/2\right)$ (82)
@@
$\frac{\mathcal{L}}{2}\left(\frac{\eta}{\sigma_{t,\text{pert}}}\right)^{2}\mathbb{E}[\|g_{\text{pert}}\|^{2}]$ (84)
$\leq\frac{\mathcal{L}}{2}\left(\frac{\eta}{\sigma_{t,\text{pert}}}\right)^{2}\left(2\|\nabla L(\theta_{t})\|^{2}+O(\rho\mathcal{L}\|\nabla L(\theta_{t})\|)+O\left(\frac{d(\sigma^{2}+\epsilon^{2}G^{2})}{\epsilon^{2}m}\right)\right).$ (85)
Since $\sigma_{t,\text{pert}}\approx\epsilon\|\nabla L(\theta_{\text{pert}})\|$, effective $\frac{\eta}{\sigma_{t,\text{pert}}}\approx\frac{\eta}{\epsilon\|\nabla L(\theta_{\text{pert}})\|}$, but for descent guarantee with small $\eta$, the quadratic term is controlled by $\eta\leq\frac{1}{2\mathcal{L}}$, yielding:
$\mathbb{E}[L(\theta_{t+1})-L(\theta_{t})]$ $\leq-\frac{\eta}{2\epsilon}\|\nabla L(\theta_{t})\|^{2}+\rho\mathcal{L}\|\nabla L(\theta_{t})\|^{2}/2+O\left(\frac{\eta}{\epsilon}\epsilon^{2}d\mathcal{L}\right)+\frac{\mathcal{L}\eta^{2}}{2\epsilon^{2}}O(1)$ (86)
@@
$\frac{\eta}{2\epsilon}\|\nabla L(\theta_{t})\|^{2}$ $\leq L(\theta_{t})-\mathbb{E}[L(\theta_{t+1})]+\rho\mathcal{L}\|\nabla L(\theta_{t})\|^{2}/2+O(\epsilon d\mathcal{L}\eta)+O\left(\frac{\mathcal{L}\eta^{2}}{\epsilon^{2}}\right)$ (88)
@@
$\frac{\eta}{2\epsilon}\sum_{t=1}^{T}\|\nabla L(\theta_{t})\|^{2}$ $\leq L(\theta_{1})-L^{*}+\frac{\rho\mathcal{L}}{2}\sum_{t=1}^{T}\|\nabla L(\theta_{t})\|^{2}+O(T\epsilon d\mathcal{L}\eta)+O\left(T\frac{\mathcal{L}\eta^{2}}{\epsilon^{2}}\right)$ (90)
@@
using Cauchy-Schwarz for the variance term: $\sum\sqrt{V_{t}}\leq\sqrt{T\sum V_{t}}$ with $V_{t}=O\left(\frac{\mathcal{L}d(\sigma^{2}+\epsilon^{2}G^{2})}{\epsilon^{2}m}\right)$. Assuming $\rho$ small such that $\frac{\rho\mathcal{L}}{2}\sum\|\nabla\|^{2}\leq\frac{\eta}{4\epsilon}\sum\|\nabla\|^{2}$ (absorbed by choice), divide by $T$:
$\frac{1}{T}\sum_{t=1}^{T}\|\nabla L(\theta_{t})\|^{2}$ $\leq\frac{2\epsilon(L(\theta_{1})-L^{*})}{\eta T}+2\rho\mathcal{L}+O(\epsilon^{2}d\mathcal{L})+O\left(\frac{\mathcal{L}\eta}{\epsilon T}(L(\theta_{1})-L^{*})\right)$ (92)
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla L(\theta_{t})\|^{2}$ $\leq O\left(\sqrt{\frac{\mathcal{L}(L(\theta_{1})-L^{*})}{T}}\right)+2\rho\mathcal{L}$ (94)
@@
For flat minima bias: ZO introduces Hessian trace bias $O(\epsilon\sqrt{d}\operatorname{Tr}(H))$ in gradient estimates, enhanced by SAM’s $\rho$-regularization approximating $\min L+\rho\|\nabla\|$, leading to stationary points with $\mathbb{E}[\operatorname{Tr}(H(\bar{\theta}))]\leq\min\operatorname{Tr}(H^{*})+O(\rho+\epsilon\sqrt{d})$.
@@
We employ the PAC-Bayesian framework to bound the generalization error. The PAC-Bayesian theorem states that, for any prior distribution $p$ and posterior distribution $q$ (possibly dependent on $\mathcal{S}$), with probability at least $1-\kappa$ over the draw of $\mathcal{S}$:
$\mathbb{E}_{q(x)}\left[\mathbb{E}_{P(X,y)}[F(x;(X,y))]\right]\leq\mathbb{E}_{q(x)}[F(x;\mathcal{S})]+\sqrt{\frac{\mathrm{KL}(q\parallel p)+\log\left(\frac{M}{\kappa}\right)}{M-1}}.$ (96)
We set the posterior $q=\mathcal{N}(\theta_{t},\epsilon I)$ for small $\epsilon>0$, approximating the deterministic parameter $\theta_{t}$ at iteration $t$.
To handle the adaptive neighborhood, we define an effective radius $r_{t}=\rho/\sigma_{t}$, where $\sigma_{t}$ is the adaptive standard deviation computed in ZOSA. We select a prior $p=\mathcal{N}(0,\sigma_{p}^{2}I)$ from a discretized set $\mathcal{T}=\{c\exp(j/k)\mid j\in\mathbb{Z}\}$ for some constant $c>0$, ensuring coverage of possible $\sigma_{p}^{2}$ values close to $r_{t}^{2}/(1+\sqrt{\log M/k})^{2}$.
Substituting into the PAC-Bayesian bound and applying a union bound over $|\mathcal{T}|$ priors (with $|\mathcal{T}|\approx\log M$ for appropriate discretization), we have, with probability at least $1-\kappa$:
$\mathbb{E}_{x\sim q}\left[\mathbb{E}_{P(X,y)}[F(x;(X,y))]\right]$ $\leq\mathbb{E}_{x\sim q}[F(x;\mathcal{S})]$ (97)
$+\sqrt{\frac{\mathrm{KL}(q\parallel p)+\log\left(\frac{M|\mathcal{T}|}{\kappa}\right)+\tilde{O}(1)}{M-1}},$ (98)
for the optimal $p\in\mathcal{T}$.
The KL divergence for Gaussian $q$ and $p$ is:
$\mathrm{KL}(q\parallel p)=\frac{1}{2}\left[k\log\frac{\sigma_{p}^{2}}{\epsilon^{2}}+k\left(\frac{\epsilon^{2}}{\sigma_{p}^{2}}-1\right)+\frac{\|\theta_{t}\|_{2}^{2}}{\sigma_{p}^{2}}\right].$ (99)
By choosing $\sigma_{p}^{2}\approx r_{t}^{2}/(1+\sqrt{\log M/k})^{2}$ and taking $\epsilon\to 0$ (with appropriate bounding), the dominant term simplifies to:
$\mathrm{KL}(q\parallel p)\leq k\log\left(1+\frac{\|\theta_{t}\|_{2}^{2}}{r_{t}^{2}\left(1+\sqrt{\frac{\log M}{k}}\right)^{2}}\right)+\tilde{O}(1).$ (100)
Incorporating the union bound, $\log(|\mathcal{T}|M/\kappa)\leq 4\log(M/\kappa)+\tilde{O}(1)$ for suitable discretization.
The sharpness-aware objective in ZOSA minimizes the maximum loss in the adaptive $L_{2}$ neighborhood:
$\max_{\|\delta\|_{2}\leq r_{t}}F(\theta_{t}+\delta;\mathcal{S}).$ (101)
Since $\mathbb{E}_{x\sim q}[F(x;\mathcal{S})]\leq\max_{\|\delta\|_{2}\leq r_{t}}F(\theta_{t}+\delta;\mathcal{S})$ (as $q$ concentrates around $\theta_{t}$ and the neighborhood captures perturbations up to $r_{t}$ via concentration inequalities for Gaussian perturbations approximating the $L_{2}$ ball),
we refine the bound:
@@
$\mathbb{E}_{x\sim q}\left[\mathbb{E}_{P(X,y)}[F(x;(X,y))]\right]$ $\leq\max_{\|\delta\|_{2}\leq r_{t}}F(\theta_{t}+\delta;\mathcal{S})$ (102)
@@
$+\sqrt{\frac{k\log\left(1+\frac{\|\theta_{t}\|_{2}^{2}}{r_{t}^{2}\left(1+\sqrt{\frac{\log M}{k}}\right)^{2}}\right)+4\log\frac{M}{\kappa}+\tilde{O}(1)}{M-1}}.$ (103)
Assuming the loss function $F(x;(X,y))$ is convex with respect to $x$, Jensen’s inequality applies in the limit $\epsilon\to 0$:
$\mathbb{E}_{P(X,y)}[F(\theta_{t};(X,y))]$ $=\mathbb{E}_{P(X,y)}[F(\mathbb{E}_{x\sim q}[x];(X,y))]$ (104)
$\leq\mathbb{E}_{x\sim q}\left[\mathbb{E}_{P(X,y)}[F(x;(X,y))]\right].$ (105)
Combining this with the previous bound yields the final result:
$\mathbb{E}_{P(X,y)}[F(\theta_{t};(X,y))]\leq\max_{\|\delta\|_{2}\leq\rho/\sigma_{t}}F(\theta_{t}+\delta;\mathcal{S})$ (106)
$+\sqrt{\frac{k\log\left(1+\frac{\|\theta_{t}\|_{2}^{2}}{(\rho/\sigma_{t})^{2}\left(1+\sqrt{\frac{\log M}{k}}\right)^{2}}\right)+4\log\frac{M}{\kappa}+\tilde{O}(1)}{M-1}}.$ (107)

### v2

@@ A.5.1 SAM Mechanism
A.5.2 ZOSA Approximation
ZOSA operates in a zero-order setting, meaning it does not have access to exact gradients. Instead, it estimates gradients using function evaluations along random directions. The step-by-step derivation of how ZOSA simulateties SAM is shown below.
@@ A.5.2 ZOSA Approximation
    $\epsilon_{\text{sam}}\approx\frac{\rho}{\epsilon}\frac{\nabla L(\theta_{t})}{\|\nabla L(\theta_{t})\|}.$ (66)
@@ A.5.3 Conclusion for Approximate Equivalence
The proof builds on standard non-convex descent lemmas, ZO estimation bias/variance bounds, SAM perturbation approximations, and Property 3.2 ($\sigma_{t}\approx\epsilon\|\nabla L(\theta_{t};\mathcal{B}_{t})\|$).
Assume the loss function $L(\theta)$ is $\mathcal{L}$-smooth and non-convex with lower bound $L^{*}$, perturbations satisfy $\mathbb{E}[(l_{i}-L(\theta))^{2}]\leq\sigma^{2}$, $\|\nabla L(\theta;\mathcal{B})\|\leq G$, $\eta_{t}\leq\frac{m}{16d\mathcal{L}}$, $\epsilon\leq\frac{1}{\sqrt{d\mathcal{L}}}$, $\rho\leq\frac{1}{4\mathcal{L}}$.
By $\mathcal{L}$-smoothness:
@@
where $\Delta\theta=-\frac{\eta}{\sigma_{t,\text{pert}}}g_{\text{pert}}$.
Substitute:
$L(\theta_{t+1})\leq L(\theta_{t})-\frac{\eta}{\sigma_{t,\text{pert}}}\langle\nabla L(\theta_{t}),\hat{g}_{\text{pert}}\rangle+\frac{\mathcal{L}}{2}\left(\frac{\eta}{\sigma_{t,\text{pert}}}\right)^{2}\|\hat{g}_{\text{pert}}\|^{2}.$ (73)
@@
$\mathbb{E}[L(\theta_{t+1})]\leq L(\theta_{t})-\frac{\eta}{\sigma_{t,\text{pert}}}\langle\nabla L(\theta_{t}),\mathbb{E}[\hat{g}_{\text{pert}}]\rangle+\frac{\mathcal{L}}{2}\left(\frac{\eta}{\sigma_{t,\text{pert}}}\right)^{2}\mathbb{E}[\|\hat{g}_{\text{pert}}\|^{2}].$ (74)
From ZO properties at perturbed point $\theta_{\text{pert}}=\theta_{t}+\rho\frac{\hat{g}_{t}}{\sigma_{t}}$:
$\mathbb{E}[\hat{g}_{\text{pert}}]=\nabla L(\theta_{\text{pert}})+O(\epsilon^{2}),$ (75)
$\mathbb{E}[\|\hat{g}_{\text{pert}}-\mathbb{E}[\hat{g}_{\text{pert}}]\|^{2}]\leq O\left(\frac{d+\epsilon^{2}G^{2}}{\epsilon^{2}m}\right),$ (76)
$\mathbb{E}[\|\hat{g}_{\text{pert}}\|^{2}]=\|\nabla L(\theta_{\text{pert}})\|^{2}+O\left(\frac{d+\epsilon^{2}G^{2}}{\epsilon^{2}m}\right).$ (77)
By smoothness, $\|\nabla L(\theta_{\text{pert}})-\nabla L(\theta_{t})\|\leq(\rho/\epsilon)\mathcal{L}$, since $\left\|\frac{\hat{g}_{t}}{\sigma_{t}}\right\|\approx 1$ (from Property 3.2, $\sigma_{t}\approx\epsilon\|\nabla L(\theta_{t})\|$, and $\hat{g}_{t}\approx\nabla L(\theta_{t})$ up to scaling adjustment for one-sided Rademacher).
Thus:
$\langle\nabla L(\theta_{t}),\nabla L(\theta_{\text{pert}})\rangle$ $\geq\|\nabla L(\theta_{t})\|^{2}-(\rho/\epsilon)\mathcal{L}\|\nabla L(\theta_{t})\|,$ (78)
$\langle\nabla L(\theta_{t}),\mathbb{E}[\hat{g}_{\text{pert}}]\rangle$ $\geq\|\nabla L(\theta_{t})\|^{2}-(\rho/\epsilon)\mathcal{L}\|\nabla L(\theta_{t})\|+O(\epsilon^{2}\|\nabla L(\theta_{t})\|).$ (79)
For bounding, assume $\eta/\sigma_{t,\text{pert}}\leq\eta/(\epsilon\sqrt{\delta})$ for small $\delta>0$ (handling near-zero division), and $\|\nabla L(\theta_{\text{pert}})\|^{2}\leq 2\|\nabla L(\theta_{t})\|^{2}+2((\rho/\epsilon)\mathcal{L})^{2}$:
$-\frac{\eta}{\sigma_{t,\text{pert}}}\langle\nabla L(\theta_{t}),\mathbb{E}[\hat{g}_{\text{pert}}]\rangle$ $\leq-\frac{\eta}{\sigma_{t,\text{pert}}}\left(\frac{1}{2}\|\nabla L(\theta_{t})\|^{2}-(\rho/\epsilon)\mathcal{L}\|\nabla L(\theta_{t})\|^{2}/2\right)$ (80)
@@
$\frac{\mathcal{L}}{2}\left(\frac{\eta}{\sigma_{t,\text{pert}}}\right)^{2}\mathbb{E}[\|\hat{g}_{\text{pert}}\|^{2}]$ (82)
$\leq\frac{\mathcal{L}}{2}\left(\frac{\eta}{\sigma_{t,\text{pert}}}\right)^{2}\left(2\|\nabla L(\theta_{t})\|^{2}+O((\rho/\epsilon)\mathcal{L}\|\nabla L(\theta_{t})\|)+O\left(\frac{d+\epsilon^{2}G^{2}}{\epsilon^{2}m}\right)\right).$ (83)
Since $\sigma_{t,\text{pert}}\approx\epsilon\|\nabla L(\theta_{\text{pert}})\|$, effective $\frac{\eta}{\sigma_{t,\text{pert}}}\approx\frac{\eta}{\epsilon\|\nabla L(\theta_{\text{pert}})\|}$, but for a small descent guaranty $\eta$, the quadratic term is controlled by $\eta\leq\frac{1}{2\mathcal{L}}$, yielding:
$\mathbb{E}[L(\theta_{t+1})-L(\theta_{t})]$ $\leq-\frac{\eta}{2\epsilon}\|\nabla L(\theta_{t})\|^{2}+(\rho/\epsilon)\mathcal{L}\|\nabla L(\theta_{t})\|^{2}/2+O\left(\frac{\eta}{\epsilon}\epsilon^{2}d\mathcal{L}\right)+\frac{\mathcal{L}\eta^{2}}{2\epsilon^{2}}O(1)$ (84)
@@
$\frac{\eta}{2\epsilon}\|\nabla L(\theta_{t})\|^{2}$ $\leq L(\theta_{t})-\mathbb{E}[L(\theta_{t+1})]+(\rho/\epsilon)\mathcal{L}\|\nabla L(\theta_{t})\|^{2}/2+O(\epsilon d\mathcal{L}\eta)+O\left(\frac{\mathcal{L}\eta^{2}}{\epsilon^{2}}\right)$ (86)
@@
$\frac{\eta}{2\epsilon}\sum_{t=1}^{T}\|\nabla L(\theta_{t})\|^{2}$ $\leq L(\theta_{1})-L^{*}+\frac{(\rho/\epsilon)\mathcal{L}}{2}\sum_{t=1}^{T}\|\nabla L(\theta_{t})\|^{2}+O(T\epsilon d\mathcal{L}\eta)+O\left(T\frac{\mathcal{L}\eta^{2}}{\epsilon^{2}}\right)$ (88)
@@
using Cauchy-Schwarz for the variance term.
Assuming $\rho$ small such that $\frac{(\rho/\epsilon)\mathcal{L}}{2}\sum\|\nabla\|^{2}\leq\frac{\eta}{4\epsilon}\sum\|\nabla\|^{2}$ (absorbed by choice), divide by T:
$\frac{1}{T}\sum_{t=1}^{T}\|\nabla L(\theta_{t})\|^{2}$ $\leq\frac{2\epsilon(L(\theta_{1})-L^{*})}{\eta T}+2(\rho/\epsilon)\mathcal{L}+O(\epsilon^{2}d\mathcal{L})+O\left(\frac{\mathcal{L}\eta}{\epsilon T}(L(\theta_{1})-L^{*})\right)$ (90)
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla L(\theta_{t})\|^{2}$ $\leq O\left(\sqrt{\frac{\mathcal{L}(L(\theta_{1})-L^{*})}{T}}\right)+2(\rho/\epsilon)\mathcal{L}$ (92)
@@
For flat minima bias: ZO introduces Hessian trace bias $O(\epsilon\sqrt{d}\operatorname{Tr}(H))$ in gradient estimates, enhanced by SAM’s $\rho$-regularization approximating $\min L+\rho\|\nabla\|$, leading to stationary points that empirically exhibit flatter minima compared to baselines.
@@
We follow the recent flatness-aware PAC-Bayesian framework of hellstrom2024generalization; andriushchenko2024generalization; zhou2025generalization, which yields dimension-independent bounds for sharpness-aware optimizers without requiring convexity.
Let $P=\mathcal{N}(0,\lambda^{-1}I_{d})$ be a data-independent Gaussian prior with $\lambda>0$. Define the data-dependent posterior
$Q\triangleq\mathcal{N}\left(\theta_{T},\,\left(\frac{\sigma_{T}^{+}}{\rho}\right)^{2}I_{d}\right),$ (94)
where $\sigma_{T}^{+}$ is the loss standard deviation observed at the final perturbed point $\theta_{T}+\delta_{T}$ (computed exactly as in ZOSA). By construction, when the minimum is flat, $\sigma_{T}^{+}$ is large $\Rightarrow$ posterior variance is large $\Rightarrow$ KL is small.
The KL divergence admits the closed-form
$D_{\mathrm{KL}}(Q\|P)=\frac{\lambda}{2}\|\theta_{T}\|_{2}^{2}+\frac{d}{2}\left[\frac{\sigma_{T}^{+2}}{\rho^{2}}(\lambda-1)+\log\frac{\rho^{2}}{\sigma_{T}^{+2}}\right]^{+}\leq O\!\left(\lambda\|\theta_{T}\|_{2}^{2}+d\log\frac{\rho}{\sigma_{T}^{+}}\right),$ (95)
where $x]^{+}=\max(x,0)$. In flat regions (large $\sigma_{T}^{+}$), the $\log(\rho/\sigma_{T}^{+})$ term becomes negative, dramatically reducing the KL.
By the PAC-Bayes-kl theorem (catoni2007pac; hellstrom2024generalization), with probability $\geq 1-\delta$,
@@
$\mathrm{kl}\!\left(\mathbb{E}_{\theta\sim Q}[\mathcal{L}_{S}(\theta)]\,\Big\|\,\mathbb{E}_{\theta\sim Q}[\mathcal{L}_{\mathcal{D}}(\theta)]\right)\leq\frac{D_{\mathrm{KL}}(Q\|P)+\log(2\sqrt{M}/\delta)}{M-1}.$ (96)
Since ZOSA explicitly minimizes the worst-case loss in the effective $\rho/\epsilon$-ball, we have (by one-step SAM approximation)
$\mathbb{E}_{\theta\sim Q}[\mathcal{L}_{S}(\theta)]\leq\mathcal{L}_{S}(\theta_{T}+\delta_{T})\leq\mathcal{L}_{S}(\theta_{T})+\rho\cdot\hat{\sigma}_{T}^{+}/\epsilon.$ (97)
But $\hat{\sigma}_{T}^{+}\approx\epsilon\|\nabla\mathcal{L}_{S}(\theta_{T})\|_{2}$ (Proposition 3.1), so
$\mathbb{E}_{\theta\sim Q}[\mathcal{L}_{S}(\theta)]\leq\mathcal{L}_{S}(\theta_{T})+\rho\cdot\operatorname{Sharp}_{\rho/\epsilon}(\theta_{T})+o(1).$ (98)
Applying kl-inversion (dziugaite2021role, Lemma A.1) and the flatness-aware bound of zhou2025generalization, the generalization gap is bounded by
@@
$\mathcal{L}_{\mathcal{D}}(\theta_{T})\leq\mathcal{L}_{S}(\theta_{T})+\rho\cdot\operatorname{Sharp}_{\rho/\epsilon}(\theta_{T})+\tilde{O}\!\left(\sqrt{\frac{D_{\mathrm{KL}}(Q\|P)+\log(M/\delta)}{M}}\right).$ (99)
The ZO estimation error contributes an additional $\tilde{O}(\sqrt{d\epsilon^{2}/m})$ term (shu2025), yielding the stated bound.